# Creative and geometric times in physics, mathematics, logic, and philosophy

### v1

@@ VI.2 Quantum theory
Quantum mechanics is the first physical theory in which indeterminism has been widely considered at a fundamental level. Indeed, Heisenberg uncertainty principle (which would be better named “indeterminacy”’ principle) and later on violation of Bell’s inequalities are strong indications that properties of quantum systems do not exist predetermined [47]. In this, quantum physics has inspired us to develop the concept of creative time, which happens at the moment when a quantum event, e.g. a measurement, determines a physical property (i.e. a single real outcome) that was previously only potential (as encapsulated by the probability amplitudes in the quantum state) [4, 48].
@@
Creative time thus becomes associated with how, and under what circumstances the potentialities become actual, i.e. to the notorious measurement problem (also at the classical level, see [3, 4]). In fact, the concept of creative time is compatible only with these classes of interpretations that conceive a passage from the potential to the determinate [49]. In quantum theory these are the ``````` `spontaneous collapse theories” or interpretations , à la Copenhagen, where the collapse is induced by (some features of) the measurement. Other interpretations, such as the many-world, cannot accommodate the concept of creative time.
@@ VI.3 Thermodynamics
A long standing problem that would be impossible not to mention is why do we ubiquitously observe a direction of time while the microscopic laws of physics are time-reversal invariant. This is commonly explained through statistical consideration, i.e., the fact that the thermodynamic quantity entropy cannot decrease in a closed system (second law of thermodynamics) [50]. This tension between the observed asymmetry in time and the time symmetry of the underlying microphysics are exemplified by the Loschmidt Paradox and by the Zermelo objection that Boltzmann’s H-Theorem is at odds with Poincaré recurrence theorem.
However, this is not necessarily the case even at the classical level, as discussed in detail in the works of B. Drossel and G. Ellis [8, 51, 52, 53, 54].
In a similar fashion, our distinction between creative and geometrical time can help clarify this matter. If there are fundamentally indeterminate events that require creative time, then there is also a fundamental asymmetry between past and future, as encapsulated by the second law. So, while admittedly this still requires to address the measurement problem (classical or quantum), paradoxes like the ones of Loschmidt or Zermelo are solved by merely noticing that Poincaré recurrence theorem simply fails [3].
@@
We join T. Maudlin when he states that ``````` `it is a fundamental, irreducible fact about the spatio-temporal structure of the world that time passes” [55]. However, the formalization of physical theories has to a large extent expelled the real passage of time from the description of natural phenomena (see also [1]). We have shown that to understand physics we should rethink our concept of time and distinguish between a geometric time (that appears in the deterministic equations of motion) and a creative time (that happens concurrently with events of actualization of potentialities).
In physics, indeed, we have related these two conceptions of time to, respectively, determinism and indeterminism. We then showed a parallel in classical and intuitionistic mathematics. The former takes a Platonistic approach by assuming that mathematical entities (such as real numbers) are given all at once, therefore avoiding the necessity of time. On the other hand, we showed that a naturalistic interpretation of intuitionistic mathematics requires a concept of creative time. Consequently, we have advocated a tensed multivalued logic (which can be true, false or indeterminate) for scientific propositions. Finally, we have related our discussion to known philosophical theories of time, showing in particular that geometric time fits best with the worldview of B-theory, whereas creative time with A-theory. Creative time has also led us to conceive a variation to the Growing-block theory, which however overcomes fundamental criticisms by providing a naturalistic account of the difference between past, present and future. These parallels in different disciplines are summarized in Table 1.
We conclude by stating that, in our view, physics is not only about sophisticated theories and fascinating technologies, it should also allow one to tell stories about “how nature does it”. But all stories require time [19]. Hence, we would like to rephrase a famous aphorism by French writer F. Rabelais as: “science without time is but ruin of intelligibility”. [^14] The original sentence reads: ``````` `Science without consciousness is but ruin of the soul”.
@@ Acknowledgements.
We thank Barbara Drossel, George Ellis, Carl Posy, Cristian Mariani, Mauro Dorato, Tim Maudlin, Valerio Scarani, Clive Aw, Mark van Atten, Andrea Oldofredi, and Baptiste Le Bihan for useful discussions and for pointing out relevant literature. This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.

### v2

@@ VI.2 Quantum theory
Quantum mechanics is the first physical theory in which indeterminism has been widely – though not universally – regarded as fundamental. Indeed, Heisenberg uncertainty principle (which would be better named “indeterminacy”’ principle) and, later on, the violation of Bell’s inequalities provided strong indications that properties of quantum systems do not exist predeterminate [69]. In this, quantum physics has inspired us to develop the concept of creative time, which happens at the moment when a quantum event, e.g. a measurement, determines a physical property (i.e. a single real outcome) that was previously only potential (as encapsulated by the probability amplitudes in the quantum state) [26, 3].
@@
Creative time thus becomes associated with how, and under what circumstances the potentialities become actual, i.e. to the notorious measurement problem (also at the classical level, see [24, 26]). In fact, the concept of creative time is compatible only with these classes of interpretations that conceive a passage from the potential to the determinate [56]. In quantum theory these are the ``spontaneous collapse theories” or interpretations, à la Copenhagen, where the collapse is induced by (some features of) the measurement. Other interpretations – such as the many-world and Bohmian mechanics – cannot accommodate the concept of creative time.
It should be noted that, while indeterminacy in classical and quantum theory leads to some common features, it also presents fundamental differences (see Ref. [27] for a detailed discussion). In quantum mechanics, certain physical quantities are incompatible (represented by noncommuting operators), and their indeterminacies are linked through Heisenberg’s uncertainty relations—for example, position and momentum. In contrast, while classical physics can be supplemented with noncontextual hidden variables (such as the real numbers used in the theory), any completion of quantum physics would require contextual and nonlocal additional variables. In philosophical terms, classical indeterminacy is considered shallow, whereas quantum indeterminacy is deep [63, 27].
@@ VI.3 Thermodynamics
A long standing problem that would be impossible not to mention is why do we ubiquitously observe a direction of time while the microscopic laws of physics are time-reversal invariant. This is commonly explained through statistical consideration, i.e., the fact that the thermodynamic quantity entropy cannot decrease in a closed system (second law of thermodynamics) [1]. This tension between the observed asymmetry in time and the time symmetry of the underlying microphysics are exemplified by the Loschmidt Paradox and by the Zermelo objection that Boltzmann’s H-Theorem is at odds with Poincaré recurrence theorem. However, this is not necessarily the case even at the classical level, as discussed in detail in the works of B. Drossel and G. Ellis [33, 34, 35, 37, 39].
In a similar fashion, our distinction between creative and geometrical time can help clarify this matter. If there are fundamentally indeterminate events that require creative time, then there is also a fundamental asymmetry between past and future, as encapsulated by the second law. So, while admittedly this still requires to address the measurement problem (classical or quantum), paradoxes like the ones of Loschmidt or Zermelo are solved by merely noticing that Poincaré recurrence theorem simply fails [24]. The consequences of indeterminacy – and the role of creative time – in statistical mechanics, arguably a domain where we possess a physical explanation of time’s direction, require further analysis and are the subject of ongoing work.
@@
We join T. Maudlin when he states that ``it is a fundamental, irreducible fact about the spatio-temporal structure of the world that time passes” [58], or, in our jargon, that time occurs. However, the formalization of physical theories has to a large extent expelled the real passage of time from the description of natural phenomena (see also [82]). We have shown that to understand physics in indeterministic terms – and we argued that there is strong conceptual and empirical ground to do so – we should rethink our concept of time and distinguish between a geometric time (that appears in the deterministic equations of motion) and a creative time (that processes concurrently with the events of actualization of potentialities).
In physics, indeed, we have tried to relate these two conceptions of time to, respectively, determinism and indeterminism. We then showed a parallel in classical and intuitionistic mathematics. The former takes a Platonistic approach by assuming that mathematical entities (such as real numbers) are given all at once, therefore avoiding the necessity of creative time. On the other hand, we showed that a naturalistic interpretation of intuitionistic mathematics requires a concept of creative time [8].We have therefore advocated a set of desiderata for a logical framework that seem to best capture indeterminism and creative time: namely, a logic that incorporates temporality (either syntactically or purely semantically) and that, by adopting an open-future semantics, rejects the law of the excluded middle. Finally, we have related our discussion to some philosophical theories of time, showing in particular that geometric time fits best with the worldview of B-theories, whereas creative time with A-theories. Creative time has also led us to conceive a variation to the Growing-block theory, which however overcomes fundamental criticisms by providing a naturalistic account of the difference between past, present, and future. Notably, we showed that to be consistent with relativity, there are many local presents that cannot be collected into a global present. These parallels in different disciplines are summarized in Table 1.
We conclude by stating that, in our view, physics is not only about sophisticated theories and fascinating technologies, it should also allow one to tell stories about “how nature does it”. But all stories require time [30]. Hence, we would like to rephrase a famous aphorism by French writer F. Rabelais as: “science without time is but ruin of intelligibility”. [^23] The original sentence reads: ``Science without consciousness is but ruin of the soul”.
@@ Acknowledgements.
We thank Clive Aw, Bruno Bentzen, Jeremy Butterfield, Mauro Dorato, Barbara Drossel, George Ellis, Michael Esfeld, Baptiste Le Bihan, Cristian Mariani, Tim Maudlin, Andrea Oldofredi, Carl Posy, Valerio Scarani, Christian Wüthrich and Mark van Atten, for useful discussions and for pointing out relevant literature. This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.